OpenStudy (anonymous):

list the three complex roots of -1 by rewriting -1 in trigonometric form and applying DeMoivre's theorem. You can leave the answer in trigonometric form.

OpenStudy (anonymous):

ready?

OpenStudy (unklerhaukus):

\[-1=e^{i\pi}\]

OpenStudy (anonymous):

omg someone finally answered praise the lord

OpenStudy (anonymous):

first we can find them without any demoivre by writing \[x^3=-1\] \[x^3+1=0\] \[(x+1)(x^2-x+1)=0\] so solutions are \(x=-1\) for the first factor, and quadratic equation finds the other two as \(x=\frac{1}{2}\pm \frac{\sqrt{3}}{2}i\)

OpenStudy (anonymous):

as for demoivre we can do that as well

OpenStudy (anonymous):

first write \(-1\) in trig form, which is kind of silly since it is a real number but that is ok in trig form it is \[-1=\cos(\pi)+i\sin(\pi)\]

OpenStudy (anonymous):



OpenStudy (anonymous):

hope that part is clear, the angle is obviously \(\pi\) from the picture

OpenStudy (anonymous):

yeah its clear lol

OpenStudy (anonymous):

now we want the cube root, and we find that by taking the real cube root of 1, which is 1, and dividing the angle by 3 (3 because it is the cube root) so the other root is \[\cos(\frac{\pi}{3})+i\sin(\frac{\pi}{3})\]

OpenStudy (anonymous):

this is the one that is \(\frac{1}{2}+\frac{\sqrt{3}}{2}i\) in standard form

OpenStudy (anonymous):

then you have a choice. easy way is to think "the other one is the conjugate, so it must be \(\frac{1}{2}-\frac{\sqrt{3}}{2}i\) "

OpenStudy (anonymous):

so can cos(pi/2)+isin(pi/2) because it's the square root of one?

OpenStudy (anonymous):

or you can go around the circle again since \[\cos(\pi)+i\sin(\pi)=\cos(3\pi)+i\sin(3\pi)\] divide by 3 again and get \[\cos(\pi)+i\sin(\pi)\] which is fairly silly because this is -1 which is what you started with

OpenStudy (anonymous):

and finally go around the circle one more time and get \[\cos(5\pi)+i\sin(5\pi)\] divide the angle by 3 and get \[\cos(\frac{5\pi}{3})+i\sin(\frac{5\pi}{3})\] which gives \[\frac{1}{2}-\frac{\sqrt{3}}{2}i\] as promised

OpenStudy (anonymous):

now to answer you question, no the square roots of 1 are 1 and -1

OpenStudy (anonymous):

ah true I totally forgot we are looking for -1 lol. .

OpenStudy (anonymous):

btw the snap way to do this is to locate one root and then the others will be evenly spaced around the unit circle

OpenStudy (anonymous):

for example to solve \(x^4=1\) the solutions are \(x=1,i,-1,-i\)

OpenStudy (anonymous):

and to solve \(x^6=1\) the solutions are \[1,\frac{1}{2}+\frac{\sqrt{3}}{2}i, -\frac{1}{2}+\frac{\sqrt{3}}{2}i,-1,-\frac{1}{2}+\frac{\sqrt{3}}{2}i,\frac{1}{2}-\frac{\sqrt{3}}{2}i\]

OpenStudy (anonymous):